MOLECULAR SPECTROSCOPY IN HELIUM DROPLETS

\underline{ANDREY F. VILESOV}, Department of Chemistry, University of Southern California, Los

Angeles, CA 90089, USA.

The spectroscopy of embedded molecules in the interior of large helium droplets provides a unique method for studying a finite sized quantum fluid object, and also demonstrates a very gentle and extremely cold matrix for the high resolution spectroscopy of the molecules and their complexes. Recent developments in these two emerging fields of research will be reviewed. Challenges and possible future directions will be outlined.